(-2,0) (4,0) what is the slope?
Found 2 solutions by uday1435, solver91311:
Answer by uday1435(57)

Slope = (y2 – y1)/ (x2 – x1) = (0 – 0)/ {(4 – (-2)} = 0/6 = 0
So the slope is 0

Answer by solver91311(24713)



Use the formula for the slope:



where and are the coordinates of your given points.

Plug in the numbers and do the arithmetic.


John